For what values of k will kx^2+kx+2=0 have two equal real roots?
Answer by stanbon(75887)

For what values of k will kx^2+kx+2=0 have two equal real roots?
-------------------
The discriminant must be equal to zero.
k^2-4*k*2 0
k^2-8k = 0
k(k-8) = 0
----------------
k=0; k=8
If k is either of these values, the quadratic will
have two equal real roots.
========================
Cheers,
Stan H.